# Scene4U: Hierarchical Layered 3D Scene Reconstruction from Single Panoramic Image for Your Immerse Exploration

### v1

@@ Abstract
The reconstruction of immersive and realistic 3D scenes holds significant practical importance in various fields of computer vision and computer graphics. Typically, immersive and realistic scenes should be free from obstructions by dynamic objects, maintain global texture consistency, and allow for unrestricted exploration. The current mainstream methods for image-driven scene construction involves iteratively refining the initial image using a moving virtual camera to generate the scene. However, previous methods struggle with visual discontinuities due to global texture inconsistencies under varying camera poses, and they frequently exhibit scene voids caused by foreground-background occlusions. To this end, we propose a novel layered 3D scene reconstruction framework from panoramic image, named Scene4U. Specifically, Scene4U integrates an open-vocabulary segmentation model with a large language model to decompose a real panorama into multiple layers. Then, we employs a layered repair module based on diffusion model to restore occluded regions using visual cues and depth information, generating a hierarchical representation of the scene. The multi-layer panorama is then initialized as a 3D Gaussian Splatting representation, followed by layered optimization, which ultimately produces an immersive 3D scene with semantic and structural consistency that supports free exploration. Scene4U outperforms state-of-the-art method, improving by 24.24% in LPIPS and 24.40% in BRISQUE, while also achieving the fastest training speed. Additionally, to demonstrate the robustness of Scene4U and allow users to experience immersive scenes from various landmarks, we build WorldVista3D dataset for 3D scene reconstruction, which contains panoramic images of globally renowned sites. The implementation code and dataset will be released at https://github.com/LongHZ140516/Scene4U
@@ 1 Introduction
To this end, we introduce a novel framework named Scene4U, which employs a multi-layer 3D scene reconstruction from panoramic images to produce highly realistic and coherent immersive experience scenes. The method comprises three primary stages:
@@
  • We propose Scene4U, a layered scene reconstruction framework that generates highly consistent 3D scenes from single panoramic image. Scene4U effectively addresses texture inconsistency and occlusion challenges within scenes, enabling users to freely explore the environment. It outperforms previous state-of-the-art (SOTA) methods in visual quality and provides more realistic immersive experiences.
@@ 2.1 Image Variation Based on Diffusion Models
Currently, most of the existing methods depend on multi-view images for 3D scene generation, which infer the 3D structure of a scene from images captured from various perspectives [22, 9, 23, 57]. However, these methods cannot be directly applied to single-view scene generation. In the field of single-view scene generation, existing methods such as LucidDreamer [11], RealmDreamer [40] and Text2Immersion [37] iteratively refine the scene by moving a virtual camera, gradually generating a complete 3D scene by capturing and inpainting information from different angles. However, during the iterative refinement process, these methods only utilize partial texture information of the scene, neglecting the global consistency, leading to obvious visual discontinuities in the generated 3D scene. In DreamScene360 [56], researchers used panoramic images to build 3D scenes, thereby ensuring the consistency of overall scene information. However, due to insufficient consideration of foreground-background occlusion relationships, DreamScene360 can only allow scene viewing from fixed perspectives and does not support free navigation. In contrast, we employ a layered construction and rendering approach, which removes the restrictions of fixed viewpoints and allows users to freely navigate within the scene, providing a more immersive experience.
@@ 3 Scene4U
We propose a panoramic image-driven framework for immersive 3D scene reconstruction that removes distracting elements (e.g., pedestrians and vehicles) to render a 3D scene with high visual consistency and scene integrity. The key insight is allowing users to freely explore scenes from any time. As shown in Fig. 2, Scene4U consists of three main stages. First, a text-prompt-driven diffusion model generates a target panoramic image with specific spatiotemporal properties. Next, a large language model assists in decomposing the target panoramic image into multiple layers. These decomposed layers are then processed with image inpainting and depth restoration to obtain complete layered scene information. Finally, we transform the multi-layered panoramic images into a multi-layered 3D scene by 3DGS refinement, creating an immersive scene that supports free exploration.
@@ 3.2 Layered Panorama Construction
Complex dynamic foreground objects in real-world scenes cause occlusion, making parts of the background invisible. To this end, we propose a novel layered representation method for 3D reconstruction, which eliminates the invisible background and constructs a complete
Multi-layer Segmentation. Multi-layer representation requires accurate identification of hierarchical objects in the scene. Firstly, we predefine hierarchical category representations of the scene, including sky $L_{sky}$, background $L_{background}$, foreground $L_{foreground}$ and dynamic object $L_{dynamic}$ layer, which are sequentially numbered as 3, 2, 1, and 0, respectively. Secondly, we employ a fine-grained open-vocabulary instance segmentation [8] to extract individual semantic objects from the panoramic image. We observe that the predicted instances are difficult to directly categorize into hierarchical layers based on semantic information, especially for class-agnostic masks. Therefore, we utilize LLM [1] to interactively perform semantic filtering and hierarchical classification of objects. Specifically, we input the original RGB image and corresponding segmentation masks into the LLM and use predefined hierarchical categories for prompting. The LLM then outputs the hierarchically filtered segmentation masks. Finally, we obtain multi-layer segmentation masks with spatial hierarchy, as shown in Fig. 4. Notably, we define objects that are likely disappear in the short term as the dynamic objects. Removing dynamic objects helps generate immersive scenes that are more spacious and natural. While there are challenges in the definition of categories, the segmentation results (Fig. 5) are robust to occlusion issues and geometric orderings.
@@
Figure 5: The segmentation results of the Multi-Layer Segmentation in indoor scenes.
Multi-layer Image Restoration. The occlusion presents an inevitable integrity issue in layered image for 3D scene construction. To address this, we use the FLUX-inpainting to progressively repair each layer using the hierarchical masks. Based on the given mask (e.g., the areas to be restored), we utilize image context information to render and fill in the matching pixels. Layer-by-layer repair ensures semantic consistency across layers while restoring occluded content, resulting in high-quality generated scene. For instance, the foreground mask is used to help the repair of occluded areas in the background, while the background mask, together with the foreground mask, assists in repairing of occluded sky regions. During the repair process, we use ’no objects present’ as a prompt to guide the repair model, allowing it to seamlessly transition and restore the original details of the scene when filling in the content of each layer. This process is formalized as:
$\tilde{I}_{l+1}=\mathcal{F}_{inpaint}(\tilde{I}_{l},M_{l+1}),$ (1)
where $\mathcal{F}_{inpaint}$ is the function of FLUX-inpainting, $\tilde{I}_{l}$ and $M_{l+1}$ denote the RGB panoramic image of the $l$-th layer and the mask for the $(l+1)$-th layer, respectively. The repaired image, obtained after filling the missing regions, maintains both structural and textural consistency as well as completeness.
Multi-layer Depth Estimation and Completion. To construct a complete and spatially accurate 3D scene, we perform depth estimation and depth completion on the layered images to ensure spatial consistency across them. Inspired by the 360MonoDepth [39], we project the panorama image onto 20 overlapping perspective patches and use a pretrained monocular depth model ZoeDepth [5] to calculate depth for each projection. To address the affine ambiguities in scale and displacement, we align each projection by inputting the optimized parameters into a multi-scale and spatially varying deformation field, to get a high-resolution panoramic depth map. We use the existing depth information and RGB texture to predict and complete the masked regions without depth value, mathematically expressed as follows:
$\tilde{d}_{l}=\mathcal{F}_{depth}(\tilde{I}_{l},d_{l+1},M_{l}),$ (2)
where $\mathcal{F}_{depth}$ denotes the model for depth completion. $d_{l+1}$ represents the depth completion result for the $(l+1)$-th layer, $\tilde{I}_{l}$ is the RGB panorama of the $l$-th layer, and $M_{l}$ denotes the mask of the $l$-th layer, respectively. Notably, to ensure that the spatial relationship between the depth information of the background and sky layers remains consistent with that of the foreground layer, we perform depth estimation exclusively for the foreground layer, while using depth completion [27] for the background and sky layers.
@@ 3.3 Multi-layer Panoramas to 3D Scene
$\theta_{i}=\frac{\pi i}{H},\quad\phi_{j}=\frac{2\pi j}{W},$ (3)
$\begin{pmatrix}X\\ Y\\ Z\end{pmatrix}=d_{(i,j)}\times\begin{pmatrix}cos\phi_{j}\cdot cos\theta_{i}\\ sin\phi_{j}\\ cos\phi_{j}\cdot sin\theta_{i}.\end{pmatrix}$ (4)
@@
We find that occlusion effects lead to inconsistencies in the layered scene. Therefore, we perform independent optimization for each layer during the 3DGS refinement. Due to the clear hierarchical structure of the scene, where the back layers do not occlude the front layers, starting the optimization from the back layers helps to establish foundational background information and ensures depth consistency across layers.Therefore, We extract the 3DGS of specific layers in depth order (from the back layers to the front) for optimization. In the training phase, we set the loss function as a weighted sum of $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$:
@@ 4.4 Comparisons with Other Methods
We select LucidDreamer [11] and DreamScene360 [56] as baselines for performance comparison with our proposed Scene4U. LucidDreamer takes a single image and textual prompts as input, employing a progressive inpainting strategy to generate a 360-degree scene. This method incrementally fills and expands the image content through multiple iterations to create a complete view. Since the original LucidDreamer does not support panoramic image input, we modify it to accept an input panorama as the initial scene to ensure a fair evaluation of the baseline methods. In contrast, DreamScene360 generates a 360-degree panoramic 3DGS scene based on text prompts, converting textual descriptions into a complete and high-quality 3D scene. To ensure a fair comparison, we excluded the Climate Controller module from the comparative experiments and additionally computed the metrics without removing dynamic objects.
@@ 5 Conclusion
In this work, we propose a novel and effective framework, named Scene4U, for generating highly realistic and consistent 3D scenes through a hierarchical reconstruction strategy. First, we stylize the input panorama and perform multi-layer decomposition with the assistance of an open-vocabulary segmentor and LLMs. Subsequently, we remove dynamic and foreground objects from the scene and design layered inpainting of the occluded areas based on image texture and depth information. Finally, the multi-layer panorama is initialized as a 3DGS representation and hierarchically refined to construct immersive scenes with semantic and structural consistency. Comprehensive experimental results demonstrate that Scene4U outperforms previous SOTA methods in terms of visual quality and achieves a realistic visual experience, especially after the removal of dynamic object layers. Scene4U is also competitive in efficiency. Overall, our Scene4U provides a novel solution for reconstructing multi-temporal, dynamic-object-free, globally consistent, and freely explorable immersive 3D scenes from panoramic images.

### v2

@@ Abstract
The reconstruction of immersive and realistic 3D scenes holds significant practical importance in various fields of computer vision and computer graphics. Typically, immersive and realistic scenes should be free from obstructions by dynamic objects, maintain global texture consistency, and allow for unrestricted exploration. The current mainstream methods for image-driven scene construction involves iteratively refining the initial image using a moving virtual camera to generate the scene. However, previous methods struggle with visual discontinuities due to global texture inconsistencies under varying camera poses, and they frequently exhibit scene voids caused by foreground-background occlusions. To this end, we propose a novel layered 3D scene reconstruction framework from panoramic image, named Scene4U. Specifically, Scene4U integrates an open-vocabulary segmentation model with a large language model to decompose a real panorama into multiple layers. Then, we employs a layered repair module based on diffusion model to restore occluded regions using visual cues and depth information, generating a hierarchical representation of the scene. The multi-layer panorama is then initialized as a 3D Gaussian Splatting representation, followed by layered optimization, which ultimately produces an immersive 3D scene with semantic and structural consistency that supports free exploration. Scene4U outperforms state-of-the-art method, improving by 24.24% in LPIPS and 24.40% in BRISQUE, while also achieving the fastest training speed. Additionally, to demonstrate the robustness of Scene4U and allow users to experience immersive scenes from various landmarks, we build WorldVista3D dataset for 3D scene reconstruction, which contains panoramic images of globally renowned sites. The implementation code and dataset will be made publicly available.
@@ 1 Introduction
Recent works such as WonderWorld [52] and LayerPano3D [49] have highlighted the potential of layered representations for enhancing scene understanding. Building upon these insights, we propose a novel framework named Scene4U, which leverages real-world panoramic images to perform multi-layer 3D scene reconstruction, producing highly realistic and coherent immersive experience scenes.
@@
  • We propose Scene4U, a layered scene reconstruction framework that generates highly consistent 3D scenes from a single real-world panoramic image. Scene4U effectively addresses challenges such as texture inconsistency and occlusions within complex environments, enabling users to immerse themselves in realistic scene exploration.
@@ 2.1 Image Variation Based on Diffusion Models
Currently, most of the existing methods depend on multi-view images for 3D scene generation, which infer the 3D structure of a scene from images captured from various perspectives [23, 9, 24, 60]. However, these methods cannot be directly applied to single-view scene generation. In the field of single-view scene generation, existing methods such as LucidDreamer [11], RealmDreamer [41] and Text2Immersion [38] iteratively refine the scene by moving a virtual camera, gradually generating a complete 3D scene by capturing and inpainting information from different angles. However, during the iterative refinement process, these methods only utilize partial texture information of the scene, neglecting the global consistency, leading to obvious visual discontinuities in the generated 3D scene. In DreamScene360 [59], researchers used panoramic images to build 3D scenes, thereby ensuring the consistency of overall scene information. However, due to insufficient consideration of foreground-background occlusion relationships, DreamScene360 can only allow scene viewing from fixed perspectives and does not support free navigation.
WonderWorld [52] and LayerPano3D [49] have pioneered the concept of distinguishing the spatial relationships among objects in a scene through scene layering.
Inspired by these methods, we employ a semantic-aware layered construction and rendering approach, which removes the restrictions of fixed viewpoints and allows users to navigate within the scene, providing a more immersive experience.
@@ 3 Scene4U
We propose a panoramic image-driven framework for immersive 3D scene reconstruction that removes distracting elements (e.g., pedestrians and vehicles) to render a 3D scene with high visual consistency and scene integrity. The key insight is allowing users to explore scenes from any time. As shown in Fig. 2, Scene4U consists of three main stages. First, a text-driven diffusion model generates a target panoramic image with specific spatiotemporal properties. Next, a large language model assists in decomposing the target panoramic image into multiple layers. These decomposed layers are then processed with image inpainting and depth restoration to obtain complete layered scene information. Finally, we transform the multi-layered panoramic images into a multi-layered 3D scene by 3DGS refinement, creating an immersive scene.
@@ 3.2 Layered Panorama Construction
Complex dynamic foreground objects in real-world scenes cause occlusion, making parts of the background invisible. To this end, we propose a novel layered representation method for 3D reconstruction, which eliminates the invisible background and constructs a complete 3D scene by dividing it into multiple visual layers.
Semantic-Aware Layer Parsing. Multi-layer representation requires accurate identification of hierarchical objects in the scene. LayerPano3D[49] innovatively proposes scene layering method using instance segmentation prior. Instance segmentation methods are generally categorized into close-vocabulary [16] and open-vocabulary [8] approaches. As illustrated in Fig. 4, the former is often constrained by a predefined set of categories, which limits its adaptability to the complexity of real-world scenarios. To address this, we propose a novel approach that effectively resolves the applicability issues of existing methods in complex environments.
Firstly, we predefine hierarchical category representations of the scene, including sky $L_{sky}$, background $L_{background}$, foreground $L_{foreground}$ and dynamic object $L_{dynamic}$ layer, which are sequentially numbered as 3, 2, 1, and 0, respectively. Secondly, we employ a fine-grained open-vocabulary instance segmentation [8] to extract individual semantic objects from the panoramic image. We observe that the predicted instances are difficult to directly categorize into hierarchical layers based on semantic information, especially for class-agnostic masks. Therefore, we utilize LLM [1] to interactively perform semantic filtering and hierarchical classification of objects. Specifically, we input the original RGB image and corresponding segmentation masks into the LLM and use predefined hierarchical categories for prompting. The LLM then outputs the hierarchically filtered segmentation masks. Finally, we obtain multi-layer segmentation masks with spatial hierarchy, as shown in Fig. 5. Notably, we define objects that are likely disappear in the short term as the dynamic objects. Removing dynamic objects helps generate immersive scenes that are more spacious and natural. While there are challenges in the definition of categories, the segmentation results (Fig. 5) are robust to occlusion issues and geometric orderings.
Figure 4: The segmentation results comparison of the close-vocabulary and open-vocabulary segmentation.
@@
Occlusion-Guided Layer Recovery. The occlusion presents an inevitable integrity issue in layered image for 3D scene construction. To address this, we use the FLUX-inpainting to progressively repair each layer using the hierarchical masks. Based on the given mask (e.g., the areas to be restored), we utilize image context information to render and fill in the matching pixels. Layer-by-layer repair ensures semantic consistency across layers while restoring occluded content, resulting in high-quality generated scene. For instance, the foreground mask is used to help the repair of occluded areas in the background, while the background mask, together with the foreground mask, assists in repairing of occluded sky regions. During the repair process, we use ’no objects present’ as a prompt to guide the repair model, allowing it to seamlessly transition and restore the original details of the scene when filling in the content of each layer. This process is formalized as:
$\tilde{I}_{l+1}=Inpaint_{RGB}(\tilde{I}_{l},M_{l+1}),$ (1)
where $Inpaint_{RGB}$ is the function of FLUX-inpainting, $\tilde{I}_{l}$ and $M_{l+1}$ denote the RGB panoramic image of the $l$-th layer and the mask for the $(l+1)$-th layer, respectively. The repaired image, obtained after filling the missing regions, maintains both structural and textural consistency as well as completeness.
Hierarchical Spatial Inference. To construct a complete and spatially accurate 3D scene, we perform depth estimation and depth completion on the layered images to ensure spatial consistency across them. Inspired by LayerPano3D [49], we utilize 360MonoDepth [40] to project the panorama image onto 20 overlapping perspective patches and use a pretrained monocular depth model ZoeDepth [5] to calculate depth for each projection. To address the affine ambiguities in scale and displacement, we align each projection by inputting the optimized parameters into a multi-scale and spatially varying deformation field, to get a high-resolution panoramic depth map. We use the existing depth information and RGB texture to predict and complete the masked regions without depth value, mathematically expressed as follows:
$\tilde{D}_{l}=Inpaint_{depth}(\tilde{I}_{l},D_{l+1},M_{l}),$ (2)
where $Inpaint_{depth}$ denotes the model for depth completion. $D_{l+1}$ represents the depth completion result for the $(l+1)$-th layer, $\tilde{I}_{l}$ is the RGB panorama of the $l$-th layer, and $M_{l}$ denotes the mask of the $l$-th layer, respectively. Notably, to ensure that the spatial relationship between the depth information of the background and sky layers remains consistent with that of the foreground layer, we perform depth estimation exclusively for the foreground layer, while using depth completion [28] for the background and sky layers.
@@ 3.3 Multi-layer Panoramas to 3D Scene
$\theta_{i,j}=\frac{\pi i}{H},\quad\phi_{i,j}=-\frac{2\pi j}{W}+\pi,$ (3)
$\begin{pmatrix}X\\ Y\\ Z\end{pmatrix}=d_{(i,j)}\times\begin{pmatrix}cos\phi_{i,j}\cdot cos\theta_{i,j}\\ cos\theta_{i,j}\\ sin\theta_{i,j}\cdot cos\phi_{i,j}\end{pmatrix}$ (4)
@@
We find that occlusion effects lead to inconsistencies in the layered scene. In response to the above difficulties, LayerPano3D [49] took the lead in proposing a solution by optimizing the 3DGS model layer by layer in multi-layer panoramas using foreground asset masks to achieve scene refinement. Followed by LayerPano3D [49], we perform independent optimization for each layer during the 3DGS refinement. Due to the clear hierarchical structure of the scene, where the back layers do not occlude the front layers, starting the optimization from the back layers helps to establish foundational background information and ensures depth consistency across layers.Therefore, We extract the 3DGS of specific layers in depth order (from the back layers to the front) for optimization. In the training phase, we set the loss function as a weighted sum of $\mathcal{L}_{1}$ and $\mathcal{L}_{D-SSIM}$:
@@ 4.4 Comparisons with Other Methods
We select publicly available methods, LucidDreamer [11] and DreamScene360 [59], as baselines for performance comparison with our proposed Scene4U. LucidDreamer takes a single image and textual prompts as input, employing a progressive inpainting strategy to generate a 360-degree scene. This method incrementally fills and expands the image content through multiple iterations to create a complete view. Since the original LucidDreamer does not support panoramic image input, we modify it to accept an input panorama as the initial scene to ensure a fair evaluation of the baseline methods. In contrast, DreamScene360 generates a 360-degree panoramic 3DGS scene based on text prompts, converting textual descriptions into a complete and high-quality 3D scene. To ensure a fair comparison, we excluded the Climate Controller module from the comparative experiments and additionally computed the metrics without removing dynamic objects.
@@ 5 Conclusion
In this work, we propose a novel and effective framework, named Scene4U, for generating highly realistic and consistent 3D scenes through a hierarchical reconstruction strategy. First, we stylize the input panorama and perform multi-layer decomposition with the assistance of an open-vocabulary segmentor and LLMs. Subsequently, we remove dynamic and foreground objects from the scene and design layered inpainting of the occluded areas based on image texture and depth information. Finally, the multi-layer panorama is initialized as a 3DGS representation and hierarchically refined to construct immersive scenes with semantic and structural consistency. Comprehensive experimental results demonstrate that Scene4U outperforms previous methods in terms of visual quality and achieves a realistic visual experience, especially after the removal of dynamic object layers. Scene4U is also competitive in efficiency. Overall, our Scene4U provides a novel solution for reconstructing multi-temporal, dynamic-object-free, globally consistent, and explorable immersive 3D scenes from panoramic images.